OpenStudy (medina13):

Solve each equation. Show all your work. Round your answers to four decimal places. 7^4x=10 Please help !!!!!

OpenStudy (amour.anohs):

Can You Screenshot That Equation And Show It

OpenStudy (amour.anohs):

^ �👈🏽�🏽Are You Trying To Say That Means Squared

OpenStudy (mightychondrion):

\[7^{4x}=10\] Thus, \[\log_7{10} =4x\] therefore \[(\log_7{10} )\div4=x\]

OpenStudy (mightychondrion):

\[\ln (2)+\ln (4x-1)=5\] Thus \[\ln (8x-2)=5\] Following that, \[e ^{5}=8x-2\] Furthermore :p \[(e ^{5}+2)/8=x\]

OpenStudy (mightychondrion):

x=18.8016 To 4 decimal places

OpenStudy (amour.anohs):

\[7^{4x}\]=10 (\[7^{4x}\])=(10) 4x*((7))=(10) 4x=\[\frac{ 10 }{ 7 }\] 4x=1.183295 \[\frac{ 4x }{ 4}\]=\[\frac{ 1.183295 }{ 4 }\] (Divide both sides by 4) x=0.295824

OpenStudy (amour.anohs):

Thats The Answer When You Round :0.2958